Problem is: 7 + 4XZ + 5(2xz + 2y)
Answer by math problem solving(30)

What you need to do is:
(1) Remove the parenthesis using the distributive law, a(b+c) = ab + ac.
(2) simplify the expanded expression by combining like terms.
7 + 4xz + 5(2xz + 2y)
=7 + 4xz + 5*2xz + 5*2y
=7 + 4xz + 10xz + 10y
= 7 + 14xz + 10y